# Weighted algebraic topology, II

### 05 May 2026 v1

@@ 4.3. Metrics and preorders
The asymmetry of a $\mu$-metric, overcome by symmetrisation, is also important, and can be highlighted by an associated preorder relation. Again there are two main ways: the reflective one (of interest for $\delta\mathbb{R}$, for instance) and the coreflective one (of interest for $\mu\mathbb{R}$ and $\delta_{0}\mathbb{R}$).
(a) A $\mu$-space $X$ can be equipped with the reflective preorder $x\prec_{\infty}x^{\prime}$
(61) $\begin{array}[]{c}{\rm pr}_{\infty}\colon\mu_{\infty}\mathsf{Mtr}\to{\rm p}\mathsf{Set},\qquad\;\;\;{\rm pr}_{\infty}(X,\mu)=(X,\prec_{\infty})\\[5.0pt] x\prec_{\infty}x^{\prime}\;\;\mbox{ if }\>\;\mu(x,x^{\prime})<\infty,\end{array}$
@@
This functor preserves finite products. It also preserves all colimits, being left adjoint to the following embedding of preordered sets in (flat) $\mu$-spaces
(62) $\begin{array}[]{c}{\rm mt}_{\infty}\colon{\rm p}\mathsf{Set}\to\mu_{\infty}\mathsf{Mtr},\qquad\;\;{\rm mt}_{\infty}(Y,\prec)=(Y,\mu_{\infty}),\\[5.0pt] \mu_{\infty}(y,y^{\prime})={\overline{\infty}}\;\mbox{ if }\>y\prec y^{\prime},\>\;\infty\;\text{ otherwise.}\end{array}$
For a preordered set $(Y,\prec)$, a mapping $f\colon(X,\mu)\to(Y,\mu_{\infty})$ satisfies $\mu_{\infty}(fx,fx^{\prime})\leqslant\lambda\mu(x,x^{\prime})$ (for some $\lambda\geqslant 0$) if and only if $\mu(x,x^{\prime})<\infty$ implies $\mu_{\infty}(fx,fx^{\prime})<\infty$, that is, $x\prec_{\infty}x^{\prime}$ implies $fx\prec fx^{\prime}$.
The reflective preorder is of interest for the line $\delta\mathbb{R}$, which gets the natural order $x\leqslant x^{\prime}$ (as in [G3, G4]). It is of no use for $\mu\mathbb{R}$ and $\delta_{0}\mathbb{R}$, which get the chaotic preorder.
(b) A $\mu$-space $X$ can also be equipped with the coreflective preorder $x\prec_{0}x^{\prime}$
(63) $\begin{array}[]{c}{\rm pr}_{0}\colon\mu_{\infty}\mathsf{Mtr}\to{\rm p}\mathsf{Set},\qquad\;\;{\rm pr}_{0}(X,\mu)=(X,\prec_{0}),\\[5.0pt] x\prec_{0}x^{\prime}\;\mbox{ if }\>\mu(x^{\prime},x)\leqslant 0.\end{array}$
This functor preserves all limits: in fact it is right adjoint to a different embedding
(64) $\begin{array}[]{c}{\rm mt}_{0}\colon{\rm p}\mathsf{Set}\to\delta\mathsf{Mtr}\subset\mu_{\infty}\mathsf{Mtr},\qquad\;\;{\rm mt}_{0}(Y,\prec)=(Y,\delta_{0}),\\[5.0pt] \delta_{0}(y,y^{\prime})=0\;\mbox{ if }\>y^{\prime}\prec y,\infty\;\text{ otherwise.}\end{array}$
For a preordered set $(Y,\prec)$, a mapping $f\colon(Y,\delta_{0})\to(X,\mu)$ satisfies $\mu(fy,fy^{\prime})\leqslant\lambda\delta_{0}(y,y^{\prime})$ (for some $\lambda\geqslant 0$) if and only if $\delta_{0}(y,y^{\prime})=0$ implies $\mu(fy,fy^{\prime})=0$, that is, $y^{\prime}\prec y$ implies $fy^{\prime}\prec_{0}fy$.
For instance, the structures $\mu\mathbb{R}$ and $\delta_{0}\mathbb{R}=(\mu\mathbb{R})^{+}$ get the natural order $x\leqslant x^{\prime}$, while $\delta\mathbb{R}$ gets the discrete one.
(c) In a $\mu$-space $X$, the relation $x\prec_{0}x^{\prime}$ implies $x^{\prime}\prec_{\infty}x$ (note the reversion). In the symmetric case both preorders are equivalence relations.
4.4. Topologies and preorders on relative metric lines
@@
(65) $\begin{array}[]{cccccc}(X,\mu)&{\rm tp}^{+}X&\hat{\mu}(x,x^{\prime})&x\prec_{\infty}x^{\prime}&\check{\mu}(x,x^{\prime})&x\prec_{0}x^{\prime}\\[5.0pt] \delta\mathbb{R}&[x_{0},x_{0}+\varepsilon[&|x-x^{\prime}|&x\leqslant x^{\prime}&discrete&discrete\\[5.0pt] \mu\mathbb{R},\delta_{0}\mathbb{R}&[{\overline{\infty}},x_{0}+\varepsilon[&chaotic&chaotic&|x-x^{\prime}|&x\leqslant x^{\prime}\end{array}$
@@
(66) $\begin{array}[]{clr}{}^{\,{\scriptstyle{\wedge}}\,}\colon\mu\mathsf{Mtr}\to\;!\mu\mathsf{Mtr},&\;{\rm pr}_{\infty}\colon\mu_{\infty}\mathsf{Mtr}\to{\rm p}\mathsf{Set}&\text{(reflective),}\\[5.0pt] {}^{\,{\scriptstyle{\vee}}\,}\colon\mu\mathsf{Mtr}\to\;!\mu\mathsf{Mtr},&\;\,{\rm pr}_{0}\colon\mu_{\infty}\mathsf{Mtr}\to{\rm p}\mathsf{Set}\qquad&\text{(coreflective).}\end{array}$
We recall that the coreflective metric (and topology) of a $\mu$-space is always finer than the reflective one, as remarked in 4.2(b). Moreover $x\prec_{0}x^{\prime}$ implies $x^{\prime}\prec_{\infty}x$.
@@ 4.5. Reflective and coreflective analyses of maps
The Lipschitz maps between $\mu$-spaces can be studied using the previous tools.
(a) By reflective tools. The $\mu_{\infty}$-maps $f\colon\delta\mathbb{R}\to\delta\mathbb{R}$ can be characterised as the mappings $\mathbb{R}\to\mathbb{R}$ which are
@@
(b) By coreflective tools. The $\mu_{\infty}$-maps $f\colon\delta_{0}\mathbb{R}\to\delta_{0}\mathbb{R}$ are characterised in the same way, using now the coreflective metric and coreflective preorder of $\delta_{0}\mathbb{R}$, which are the usual ones on the line.
@@
(c) We have already seen that all Lipschitz maps $\mu\mathbb{R}\to\mu\mathbb{R}$ are affine (in 2.6(a)), and the Lipschitz maps $\mu\mathbb{R}\to\delta_{0}\mathbb{R}$ are constant (in 2.6(b)).
@@ 4.6. Another example
(a) A mapping $f\colon\delta\mathbb{R}\to\mu\mathbb{R}$ is $\lambda$-Lipschitz if and only if
@@
a form of upper-right uniform continuity condition: for every $\varepsilon>0$, if $x<x_{0}+\varepsilon$ then $f(x)\leqslant f(x_{0})+\lambda\varepsilon$. Thus, the function $f$ is $\lambda$-Lipschitz for the euclidean metric on each subinterval where it is increasing, but can have any ‘downward jump’, as suggested in the following figure and analysed below
(70) ${p}$${x}$${y}\textstyle{{\scriptscriptstyle\bullet}}\textstyle{{\scriptscriptstyle\bullet}}$
(b) The function $f$ satisfies (69) if and only if its graph-function $\varphi(x)=(x,f(x))$ takes the natural order of the line to the (partial) order $(x,y)\leqslant_{\lambda}(x^{\prime},y^{\prime})$ of the plane defined by the relations
(71) $x\leqslant x^{\prime},\qquad\;\;y^{\prime}\leqslant y+\lambda(x^{\prime}-x).$
This order is invariant up to translations; its cone of the future at a point $p$ is represented in (70), for $\lambda=1$.
(c) For a mapping $f\colon\delta\mathbb{R}\to\delta_{0}\mathbb{R}$ the $\lambda$-Lipschitz condition still amounts to condition (69).
@@ 4.7. Lemma (Chaotic symmetrisations)
(a) The reflective symmetrisation of an affordable non-positive $\mu$-metric space $X$ is a chaotic $\mu$-space $\mu_{c}|X|$.
(b) In particular, $(\mu\mathbb{R})^{\,{\scriptstyle{\wedge}}\,}=\mu_{c}\mathbb{R},$ the chaotic $\mu$-line.
@@
In fact, assuming that $\mu(z,z^{\prime})<0$ for a pair of points, we have for all $x,x^{\prime}$ in $X$:
$\hat{\mu}(x,x^{\prime})\leqslant\mu(x,z)+(2k-1)\vartheta+\mu(z^{\prime},x^{\prime}),$
where $\vartheta=\mu(z,z^{\prime}){\,{\scriptstyle{\wedge}}\,}\mu(z^{\prime},z)<0$ and $k$ is any integer >0>0. (We are using the step-path ${\underline{x}}=(x,z,z^{\prime},...,z,z^{\prime},x^{\prime})$, where the pair $(z,z^{\prime})$ is repeated $k$ times.) ∎
@@ 4.8. Linear metrics and topology
As remarked since the Introduction, relative metrics have a complex relationship to topologies. We have so far considered four topologies on a $\mu$-metric space $X$, namely $X^{\,{\scriptstyle{\vee}}\,}\geqslant{\rm tp}^{+}X,{\rm tp}^{-}X\geqslant X^{\,{\scriptstyle{\wedge}}\,}$, but there are cases of interest where the ‘space’ already bears a natural topology, even finer than $X^{\,{\scriptstyle{\vee}}\,}$.
(a) Extending the case of $\mu\mathbb{R}$, let us consider a topological space $X$ equipped with the linear $\mu$-metric $\mu=\Delta\varphi$ associated to a continuous potential $\varphi\colon X\to\mathbb{R}$, as in (39).
The topology of $X$ is (weakly) finer than the coreflective topology ${\rm tp}^{\,{\scriptstyle{\vee}}\,}(X,\mu)$, defined by the coreflective metric $\check{\mu}=\mu{\,{\scriptstyle{\vee}}\,}\mu^{{\rm op}}$: in fact, every open ball of the latter
(72) $B_{\varphi}(x_{0},\varepsilon)=\{x\in X\,|\,|\varphi(x)-\varphi(x_{0})|<\varepsilon\}\qquad\qquad(\varepsilon>0),$
is also open in $X$, as the $\varphi$-preimage of an open interval of $\mathbb{R}$.
These topologies coincide for $\mu\mathbb{R}$ and many related cases, but the original topology can be strictly finer than the coreflective one: for instance, if $\varphi$ is constant, $\mu=\check{\mu}=\hat{\mu}$ is the chaotic $\delta$-metric, constant at 0, independently of the topology of $X$.
(b) We also note that, in the previous case, the reflective topology ${\rm tp}^{\,{\scriptstyle{\wedge}}\,}(X,\mu)$ is always chaotic:
- if $\varphi$ is constant, we already know that $\hat{\mu}$ is the chaotic $\delta$-metric,
- otherwise, $\mu$ is an affordable non-positive $\mu$-metric, and $\hat{\mu}$ is the chaotic $\mu$-metric (constant at $\overline{\infty}$}), by Lemma 4.7.
5. Paths in relative metric spaces
As claimed in the Introduction, studying paths in a $\mu$-space $X$ is likely the best way to explore its topological information and ‘direction’. We can now make precise many topics loosely described in the elementary example of 0.1.
@@ 5. Paths in relative metric spaces
$X$ always denotes a $\mu$-space.
@@ 5.1. Lipschitz paths
(a) Paths (and homotopies) in $\mu$-spaces will be based on the standard $\delta$-interval $\delta\mathbb{I}$. The positivity of its metric will play an important role.
A Lipschitz path in a $\mu$-space $X$ will be a Lipschitz map $a\colon\delta\mathbb{I}\to X$. This means that there exists a (finite) $\lambda\geqslant 0$ such that
(73) $\mu_{X}(a(t),a(t^{\prime}))\leqslant\lambda(t^{\prime}-t),\qquad\mbox{for }t\leqslant t^{\prime},$
@@
(b) We note that:
(i) the constant loop $e_{x}$ at any point $x\in X$ is Lipschitz, since $\mu(x,x)$ is 0 or $\overline{\infty}$},
@@
(iii) Lipschitz paths are closed under (partial) reparametrisation, by any $\mu_{\infty}$-map $\rho\colon\delta\mathbb{I}\to\delta\mathbb{I}$; in other words, if $a\colon\delta\mathbb{I}\to X$ is Lipschitz, so is obviously $a\rho$.
(c) A Lipschitz path $\delta\mathbb{I}\to\delta\mathbb{R}$ is any increasing mapping which is Lipschitz for the euclidean metrics, as proved in 4.5(a) (for maps $\delta\mathbb{R}\to\delta\mathbb{R}$). The same applies to the reparametrisations $\rho\colon\delta\mathbb{I}\to\delta\mathbb{I}$ considered above. We also recall that a continuous endomap of $\mathbb{I}$ need not be Lipschitz: the square root is not.
(d) For a mapping $a\colon\delta\mathbb{I}\to\mu\mathbb{R}$ the $\lambda$-Lipschitz condition amounts to
@@
and has been described in 4.6 (for maps $\delta\mathbb{R}\to\mu\mathbb{R}$).

### 13 May 2026 v2

@@ 4.3. Metrics and preorders
The asymmetry of a $\rho$-metric, overcome by symmetrisation, is also important, and can be highlighted by an associated preorder relation. Again there are two main ways: the reflective one (of interest for $\delta\mathbb{R}$, for instance) and the coreflective one (of interest for $\rho\mathbb{R}$ and $\delta_{0}\mathbb{R}$).
(a) A $\rho$-space $X$ can be equipped with the reflective preorder $x\prec_{\infty}x^{\prime}$
(56) $\begin{array}[]{c}{\rm pr}_{\infty}\colon\rho_{\infty}\mathsf{Mtr}\to{\rm p}\mathsf{Set},\qquad\;\;\;{\rm pr}_{\infty}(X,\rho)=(X,\prec_{\infty})\\[5.0pt] x\prec_{\infty}x^{\prime}\;\;\mbox{ if }\>\;\rho(x,x^{\prime})<\infty,\end{array}$
@@
This functor preserves finite products. It also preserves all colimits, being left adjoint to the following embedding of preordered sets in (flat) $\rho$-spaces:
(57) $\begin{array}[]{c}{\rm mt}_{\infty}\colon{\rm p}\mathsf{Set}\to\rho_{\infty}\mathsf{Mtr},\qquad\;\;{\rm mt}_{\infty}(Y,\prec)=(Y,\rho_{\infty}),\\[5.0pt] \rho_{\infty}(y,y^{\prime})={\overline{\infty}}\;\mbox{ if }\>y\prec y^{\prime},\>\;\infty\;\text{ otherwise.}\end{array}$
For a preordered set $(Y,\prec)$, a mapping $f\colon(X,\rho)\to(Y,\rho_{\infty})$ satisfies $\rho_{\infty}(fx,fx^{\prime})\leqslant\lambda\rho(x,x^{\prime})$ (for some $\lambda\geqslant 0$) if and only if $\rho(x,x^{\prime})<\infty$ implies $\rho_{\infty}(fx,fx^{\prime})<\infty$, that is, $x\prec_{\infty}x^{\prime}$ implies $fx\prec fx^{\prime}$.
The reflective preorder is of interest for the line $\delta\mathbb{R}$, which gets the natural order $x\leqslant x^{\prime}$ (as in [G3, G4]). It is of no use for $\rho\mathbb{R}$ and $\delta_{0}\mathbb{R}$, which get the chaotic preorder.
(b) A $\rho$-space $X$ can also be equipped with the coreflective preorder $x\prec_{0}x^{\prime}$
(58) $\begin{array}[]{c}{\rm pr}_{0}\colon\rho_{\infty}\mathsf{Mtr}\to{\rm p}\mathsf{Set},\qquad\;\;{\rm pr}_{0}(X,\rho)=(X,\prec_{0}),\\[5.0pt] x\prec_{0}x^{\prime}\;\mbox{ if }\>\rho(x^{\prime},x)\leqslant 0,\end{array}$
saying that the backward transition, from $x^{\prime}$ to $x$, gives a gain $\gamma(x^{\prime},x)\geqslant 0$.
This functor preserves all limits; in fact it is right adjoint to a different embedding
(59) $\begin{array}[]{c}{\rm mt}_{0}\colon{\rm p}\mathsf{Set}\to\delta\mathsf{Mtr}\subset\rho_{\infty}\mathsf{Mtr},\qquad\;\;{\rm mt}_{0}(Y,\prec)=(Y,\delta_{0}),\\[5.0pt] \delta_{0}(y,y^{\prime})=0\;\mbox{ if }\>y^{\prime}\prec y,\infty\;\text{ otherwise.}\end{array}$
For a preordered set $(Y,\prec)$, a mapping $f\colon(Y,\delta_{0})\to(X,\rho)$ satisfies $\rho(fy,fy^{\prime})\leqslant\lambda\delta_{0}(y,y^{\prime})$ (for some $\lambda\geqslant 0$) if and only if $\delta_{0}(y,y^{\prime})=0$ implies $\rho(fy,fy^{\prime})=0$, that is, $y^{\prime}\prec y$ implies $fy^{\prime}\prec_{0}fy$.
For instance, the structures $\rho\mathbb{R}$ and $\delta_{0}\mathbb{R}=(\rho\mathbb{R})^{+}$ get the natural order $x\leqslant x^{\prime}$, while $\delta\mathbb{R}$ gets the discrete one.
(c) In a $\rho$-space $X$, the relation $x\prec_{0}x^{\prime}$ implies $x^{\prime}\prec_{\infty}x$ (note the reversion). In the symmetric case both preorders are equivalence relations.
4.4. Topologies and preorders on real metric lines
@@
(60) $\begin{array}[]{cccccc}(X,\rho)&{\rm tp}^{+}X&\hat{\rho}(x,x^{\prime})&x\prec_{\infty}x^{\prime}&\check{\rho}(x,x^{\prime})&x\prec_{0}x^{\prime}\\[5.0pt] \delta\mathbb{R}&[x_{0},x_{0}+\varepsilon[&|x-x^{\prime}|&x\leqslant x^{\prime}&discrete&discrete\\[5.0pt] \rho\mathbb{R},\delta_{0}\mathbb{R}&[{\overline{\infty}},x_{0}+\varepsilon[&chaotic&chaotic&|x-x^{\prime}|&x\leqslant x^{\prime}\end{array}$
@@
(61) $\begin{array}[]{clr}{}^{\,{\scriptstyle{\wedge}}\,}\colon\rho\mathsf{Mtr}\to\;!\rho\mathsf{Mtr},&\;{\rm pr}_{\infty}\colon\rho_{\infty}\mathsf{Mtr}\to{\rm p}\mathsf{Set}&\text{(reflective),}\\[5.0pt] {}^{\,{\scriptstyle{\vee}}\,}\colon\rho\mathsf{Mtr}\to\;!\rho\mathsf{Mtr},&\;\,{\rm pr}_{0}\colon\rho_{\infty}\mathsf{Mtr}\to{\rm p}\mathsf{Set}\qquad&\text{(coreflective).}\end{array}$
We recall that the coreflective metric (and topology) of a $\rho$-space is always finer than the reflective one, as remarked in 4.2(b). Moreover $x\prec_{0}x^{\prime}$ implies $x^{\prime}\prec_{\infty}x$.
@@ 4.5. Reflective and coreflective analyses of maps
The Lipschitz maps between $\rho$-spaces can be studied using the previous tools.
(a) By reflective tools. The $\rho_{\infty}$-maps $f\colon\delta\mathbb{R}\to\delta\mathbb{R}$ can be characterised as the mappings $\mathbb{R}\to\mathbb{R}$ which are
@@
(b) By coreflective tools. The $\rho_{\infty}$-maps $f\colon\delta_{0}\mathbb{R}\to\delta_{0}\mathbb{R}$ are characterised in the same way, using now the coreflective metric and coreflective preorder of $\delta_{0}\mathbb{R}$, which are the usual ones on the line.
@@
(c) We have already seen that all Lipschitz maps $\rho\mathbb{R}\to\rho\mathbb{R}$ are affine (in 2.6(a)), and the Lipschitz maps $\rho\mathbb{R}\to\delta_{0}\mathbb{R}$ are constant (in 2.6(b)).
@@ 4.6. Another example
(a) A mapping $f\colon\delta\mathbb{R}\to\rho\mathbb{R}$ is $\lambda$-Lipschitz if and only if
@@
a form of upper-right uniform continuity condition: for every $\varepsilon>0$, if $x<x_{0}+\varepsilon$ then $f(x)\leqslant f(x_{0})+\lambda\varepsilon$. Thus, the function $f$ is $\lambda$-Lipschitz for the euclidean metric on each subinterval where it is increasing, but can have any ‘downward jump’.
(b) For a mapping $f\colon\delta\mathbb{R}\to\delta_{0}\mathbb{R}$ the $\lambda$-Lipschitz condition still amounts to condition (64).
@@ 4.7. Lemma (Chaotic symmetrisations)
(a) The reflective symmetrisation of an affordable non-positive $\rho$-metric space $X$ is a chaotic $\rho$-space $\rho_{C}|X|$.
(b) In particular, $(\rho\mathbb{R})^{\,{\scriptstyle{\wedge}}\,}=\rho_{C}\mathbb{R},$ the chaotic $\rho$-line.
@@
In fact, assuming that $\rho(z,z^{\prime})<0$ for a pair of points, we have for all $x,x^{\prime}$ in $X$
$\hat{\rho}(x,x^{\prime})\leqslant\rho(x,z)+(2k-1)\zeta+\rho(z^{\prime},x^{\prime}),$
where $\zeta=\rho(z,z^{\prime}){\,{\scriptstyle{\wedge}}\,}\rho(z^{\prime},z)<0$ and $k$ is any integer >0>0. (We are using the step-path ${\underline{x}}=(x,z,z^{\prime},...,z,z^{\prime},x^{\prime})$, where the pair $(z,z^{\prime})$ is repeated $k$ times.) ∎
@@ 4.8. Linear metrics and topology
As remarked since the Introduction, real metrics have a complex relationship to topologies. We have so far considered four topologies on a $\rho$-metric space $X$, namely $X^{\,{\scriptstyle{\vee}}\,}\geqslant{\rm tp}^{+}X,{\rm tp}^{-}X\geqslant X^{\,{\scriptstyle{\wedge}}\,}$, but there are cases of interest where the ‘space’ already bears a natural topology, even finer than $X^{\,{\scriptstyle{\vee}}\,}$.
Extending the case of $\rho\mathbb{R}$, let us consider a topological space $X$ equipped with the linear $\rho$-metric $\rho=\Delta\Phi$ associated to a continuous potential $\Phi\colon X\to\mathbb{R}$, as in (39).
The topology of $X$ is (weakly) finer than the coreflective topology ${\rm tp}^{\,{\scriptstyle{\vee}}\,}(X,\rho)$, defined by the coreflective metric $\check{\rho}=\rho{\,{\scriptstyle{\vee}}\,}\rho^{{\rm op}}$: in fact, every open ball of the latter
(65) $B_{\Phi}(x_{0},\varepsilon)=\{x\in X\,|\,|\Phi(x)-\Phi(x_{0})|<\varepsilon\}\qquad\qquad(\varepsilon>0),$
is also open in $X$, as the $\Phi$-preimage of an open interval of $\mathbb{R}$.
These topologies coincide for $\rho\mathbb{R}$ and many related cases, but the original topology can even be strictly finer than the coreflective one: for instance, if $\Phi$ is constant, $\rho=\check{\rho}=\hat{\rho}$ is the chaotic $\delta$-metric, constant at 0, independently of the topology of $X$.
5. Paths in real valued metric spaces
As claimed in the Introduction, studying paths in a $\rho$-space $X$ is likely the best way to explore its topological information and ‘direction’. We can now make precise many topics loosely described in the elementary example of 0.2.
@@ 5. Paths in real valued metric spaces
$X$ always denotes a $\rho$-space.
@@ 5.1. Lipschitz paths
(a) Paths (and homotopies) in $\rho$-spaces will be based on the standard $\delta$-interval $\delta\mathbb{I}$. The positivity of its metric will play an important role.
A Lipschitz path in a $\rho$-space $X$ will be a Lipschitz map $a\colon\delta\mathbb{I}\to X$. This means that there exists a (finite) $\lambda\geqslant 0$ such that
(66) $\rho_{X}(a(t),a(t^{\prime}))\leqslant\lambda(t^{\prime}-t),\qquad\mbox{for }t\leqslant t^{\prime},$
@@
(b) We note that
(i) the constant loop $e_{x}$ at any point $x\in X$ is Lipschitz, since $\rho(x,x)$ is 0 or $\overline{\infty}$},
@@
(iii) Lipschitz paths are closed under (partial) reparametrisation, by any $\rho_{\infty}$-map $\varphi\colon\delta\mathbb{I}\to\delta\mathbb{I}$; in other words, if $a\colon\delta\mathbb{I}\to X$ is Lipschitz, so is obviously $a\varphi$.
(c) A Lipschitz path $\delta\mathbb{I}\to\delta\mathbb{R}$ is any increasing mapping which is Lipschitz for the euclidean metrics, as proved in 4.5(a) (for maps $\delta\mathbb{R}\to\delta\mathbb{R}$). The same applies to the reparametrisations $\varphi\colon\delta\mathbb{I}\to\delta\mathbb{I}$ considered above. We also recall that a continuous endomap of $\mathbb{I}$ need not be Lipschitz: the square root is not.
(d) For a mapping $a\colon\delta\mathbb{I}\to\rho\mathbb{R}$ the $\lambda$-Lipschitz condition amounts to
@@
and has been described in 4.6 (for maps $\delta\mathbb{R}\to\rho\mathbb{R}$).